OpenStudy (anonymous):

A 1.50 x 10^3 kg car leaves a parking lot. Thirty seconds later it is moving along a highway at 72km/hr. a. What is the car's change in momentum b. What average force does the motor produce to bring about this change in momentum.

OpenStudy (anonymous):

What is the definition of momentuM?

OpenStudy (anonymous):

Isn't it the movement of an object through intertia?

OpenStudy (anonymous):

What is the true definition?

OpenStudy (anonymous):

72 km/hr = 20 m/s change in momentum = final momentum - initial momentum = (1.5E3) (20) - 0 = 30,000 kg m/s force = mass * acceleration acceleration = (20m/s)/30s = 2/3 m/s^2 force = (1.5E3) (2/3) = 1000 N

OpenStudy (anonymous):

isn't it p72km/h−pparked=Δp

OpenStudy (anonymous):

yes, that's what i did. I just changed the unit to m/s

OpenStudy (anonymous):

i get 108000?

OpenStudy (anonymous):

1.5x10^3(72)-1.5x10^3(0) = 108000

OpenStudy (anonymous):

well, that's correct if you don't want to change to m/s. 108,000 kg km/hr = 30,000 kg m/s

OpenStudy (anonymous):

okay now I understand. Also how would we achieve b?

OpenStudy (anonymous):

find acceleration, and use F = ma

OpenStudy (anonymous):

how do we find acceleration?

OpenStudy (anonymous):

Is there a specific formula?

OpenStudy (anonymous):

acceleration = (final velocity - initial velocity)/time

OpenStudy (anonymous):

so then. Its 72-0/30

OpenStudy (anonymous):

and then p=mv

OpenStudy (anonymous):

well the unit will be different. (km/hr)/s is isn't too conventional. it better to convert to (m/s)/s

OpenStudy (anonymous):

i mean 3600kmhr/s